# leftmargin=0.15cm, rightmargin=0.15cm, font=italic, vskip=0.10cm, indentfirst=false

### v4

@@ V-A2 Sigstore’s Weaknessses
Enterprise Adoption Limitations 7 subjects
Rate Limiting Problems – T P3, P7, P14, P15
Lack of dedicated Support & Maintenance – T & P P15, P2
Not Suited for Regulated Organizations – M & O P3
Latency Concerns – T P6
Transparency Log Issues 6 subjects
Not Suitable for private Setup – T P2, P3, P6, P14, P17
Use in Air Gap Conditions – T P2, P3, P9
Efforts to Monitor Logs – P P2
Private Sigstore Instance Setup 5 subjects
Documentation – P & T P9, P12, P6
Limited Community Support – M P15
Infrastructure Requirements & Maintenance Costs – T P5, P6
Other Documentations and Usage Information IssuesP & T 3 — P1, P14, P17
Integration to Other Systems 3 subjects
Attestation Storage – T P1
Gitlab & Jenkins – T P9
Other Unsupported technologies – T P16
Offline CapabilitiesT 2 – P3, P4
Fulcio Issues 1 – P3
Timestamping Issues – T
Fulcio-OIDC Workflow – T
Software LibrariesT 1 – P7
@@
Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications. The issues reported by participants include rate limiting—where Sigstore restricts the number of signatures that can be created per unit time—along with latency concerns, where the service’s turnaround time for a large volume of signatures is problematic. Additionally, participants highlighted the lack of dedicated support and maintenance teams due to Sigstore ’s open-source nature, as well as its unsuitability for organizations in regulated sectors for the same reason. {quoting} P7 (security tools, cloud security): “We were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise [rate limit].” {quoting} P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable. And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.” The issues here were a mixture of all usability factors.
@@
Setting up Private Sigstore Instance: Sigstore’s customizability was a feature frequently utilized by participants. However, it was also associated with certain challenges. The primary issues identified included limited documentation on setup, scarce community usage information, and the infrastructure requirements for private deployments. Issues here were primarily related to participants’ need for support resources, which were macroenvironmental and social in nature. {quoting} P15 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to…deploy Fulcio, like on prem, or in their own infrastructure.”
@@
Software Libraries: The programming libraries of Sigstore were also mentioned as a challenge, particularly in cases where users wanted to integrate Sigstore’s signing capabilities directly into an application.
V-A3 Effect of Identified Problems on Sigstore components used by Subjects
@@
Among these, our data shed the most light on local deployments and the Rekor log. For local deployment, several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table IV, section V-A2). The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues. Of the eight participants who used it, three were still in the pilot phase. Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other elements was notably absent. In particular, there was very little mention of log witnessing [86], or monitoring [87] for either Fulcio or Rekor. This is remarkable, as transparency solutions require the existence of these parties to ensure log integrity and identify log misbehavior. We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
V-B RQ2: How & Why Practitioners Change Tools
V-B1 Drivers of Adoption Among Current Sigstore Users
@@
Most of our participants chose to switch to Sigstore due to poor experiences with previous tools and due to perceived advantages associated with Sigstore. We summarize these factors in Table X.
Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore. Some participants (e.g., P6) stated that their organizations have a policy of contributing to open-source communities they intend to adopt. Others noted that their organizations were relatively new compared to Sigstore—meaning that Sigstore had already existed before these organizations were established, and some of their members had contributed to Sigstore before joining their respective organizations. Additionally, some participants mentioned that key members of their organizations were directly involved in the initial development of the Sigstore project. P14 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community. There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.”
Available Sigstore Functionalities: Sigstore’s functionalities also influenced its adoption. While several participants were drawn to the transparency log (as highlighted in section V-A1). {quoting} P5 (internal security tool, cloud security): “Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, [and] they also maintain a transparency log.”
@@
Regulations and Standards: The impact of regulations and standards on adopting Sigstore was significant among our participants. Some participants who previously used other tools (2) switched to Sigstore due to security frameworks and standards recommending it as a best practice. Additionally, the remaining participants (2) noted that regulations played a partial role in their decision to adopt Sigstore.
@@
User Community & Trust Of Sigstore Creators: The large user community of Sigstore also motivated its adoption among participants. Conversely, participants P1 and P14 cited GPG’s smaller user base as a drawback. Additionally, some participants expressed trust in industry consortia such as CNCF (Cloud Native Computing Foundation) {quoting} P3 (SAAS, SSC security tool): “We do defer often at the higher-level projects to different foundations. So we would be more likely to trust something from the CNCF…[than] independent developer projects.”
GPG & Notary Issues: Issues associated with previously existing tools played a major role in practitioners’ decisions to switch to Sigstore. GPG-based signing implementations have historically been criticized for key management challenges and usability concerns. Our participants echoed these issues, citing additional factors such as low adoption rates, steep learning curves, and limited integration with modern technologies.
Notary-related issues primarily concerned customer demands, compatibility with other technologies, and infrequent updates. Key and identity management challenges were common to both GPG and Notary. Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
V-B2 Barriers and Motivators Spurring Consideration of Sigstore Among Non-Sigstore Users
TABLE V: Non Sigstore Users: Reasons for Not Adopting Sigstore. This table summarizes the reasons non-users gave for not adopting Sigstore, grouped by their current signing tool and mapped to our usability framework theme.
@@ V-B2 Barriers and Motivators Spurring Consideration of Sigstore Among Non-Sigstore Users
In this section we highlight reasons mentioned by practitioners that have stopped them (or their organization) from switching to Next-gen tools like sigstore. We also report issues experienced by these non sigstore users that have prompted a consideration (or in some cases partial adoption of these next-gen tools).
@@
Barriers to Considering Sigstore: As highlighted in Table V, most reasons for non-adoption were organizational. {quoting} P8 (internal security tool & cloud APIs, internet services): “When it comes to open source projects and open source stuff, our Organization is very skeptical. Policy and everything drives very much teams away from that sort of stuff…There’s also the ongoing risk of having a product that’s built outside continually shipping code and being integrated into this product, where it deals with sensitive material. So we do evaluate these things, but the scrutiny against code written outside of the company is extreme, and with good reason given the context of all the supply chain stuff.” Other notable issues arose from technical concerns, including privacy risks associated with Sigstore’s public instance and the inherent risk of relying on a third-party tool maintained by open-source contributors.
@@
Factors Motivating Consideration for sigstore: When practitioners were asked about issues with their current tools and potential fixes, most pointed to existing tooling problems as the primary driver for considering alternatives—such as incorporating next-gen features into internal tools or partially adopting new solutions. Although some subjects reported drawing design inspiration from next-generation workflows, none specifically mentioned Sigstore. P11 (firmware & testing software, social technology) states, “. I’m sure at some point in time it was influenced by an open source project before they adopted it and made the right tweaks internally.” Lack of transparency and integration of these tools to other platforms were also common. We summarize these factors in Table XI.
VI Discussion
VI-A Lessons for Next-Generation Tools
Next-generation signing tools offer unparalleled flexibility by integrating with external platforms, and services. However, our participants consistently highlighted integration challenges as a major obstacle.
VI-A1 Automated Cross-Platform Integrations

### v5

@@ 5.1.2 Sigstore’s Weaknessses
Technological Factors
Transparency Log Constraints 6 subjects
1. Not suitable for private setup P2, P3, P6, P14, P17
2. Use in air-gapped/offline conditions P2, P3, P9
Performance & Rate Limits 6 subjects
1. Rate limiting problems P3, P7, P14, P15
2. Latency concerns P6
Integrations & Tooling 3 subjects
1. GitLab/Jenkins P9
2. Other unsupported technologies P16
3. Attestation storage P1
Documentation (technical aspects) 6 subjects
1. Private instance setup documentation P9, P12, P15
2. Other documentation / usage issues P1, P14, P17
Private Instance: Infra & Cost 2 subjects
1. Infra requirements & maintenance cost P5, P6
Fulcio / Timestamping Workflow 1 subject
1. Timestamping issues P3
2. Fulcio–OIDC workflow P3
Software Libraries 1 subject
1. Unsupported software libraries P7
Social / Human Factors
Log Monitoring Burden 1 subject
1. Effort to monitor logs P2
Support & Maintenance 3 subjects
1. Lack of dedicated support & maintenance P15, P2, P6
Organizational Factors
Regulatory Suitability (Cross-factor) 1 subject
1. Not suited for regulated organizations — M & O P3
Macroenvironmental Factors
Community Maturity/Support 1 subject
1. Limited community support P15
@@
The primary weakness along this dimension related to Sigstore’s use in large enterprise contexts.
Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications. The issues reported by participants include rate limiting—where Sigstore restricts the number of signatures that can be created per unit time—along with latency concerns, where the service’s turnaround time for a large volume of signatures is problematic. Additionally, participants highlighted the lack of dedicated support and maintenance teams due to Sigstore’s open-source nature, as well as its unsuitability for organizations in regulated sectors for the same reason. {quoting} P7 (security tools, cloud security): “We were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise [rate limit].” {quoting} P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable. And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.”
@@
Setting up Private Sigstore Instance: Sigstore’s customizability was a feature frequently utilized by participants. However, it was also associated with certain challenges. The primary issues identified included limited documentation on setup, scarce community usage information, and the infrastructure requirements for private deployments. Issues here were primarily related to participants’ need for support resources, which were macroenvironmental and social in nature. {quoting} P15 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to…deploy Fulcio, like on premises.”
@@
Software Libraries: The programming libraries of Sigstore were also reported as a challenge, particularly when participants sought to embed signing capabilities directly into applications. They noted the absence of well-supported, easy-to-use client libraries, which forced them to rely on workarounds such as shelling out to the CLI.
5.1.3 Impact of Issues on Sigstore Component Use
@@
Among these, our data shed the most light on local deployments and the Rekor log. For local deployment, several participants cited difficulties with documentation and setup. Only two successfully deployed private instances, likely due to these challenges (Table˜4, section˜5.1.2). The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues. Of the eight participants who used it, three were still in the pilot phase. Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
Other Components Without Mention
Though the participants in the survey shared experiences with user-facing components, discussion about other elements was notably absent. In particular, there was very little mention of log witnessing [35] or monitoring [3] for either Fulcio or Rekor. This is remarkable, as transparency solutions require the existence of these parties to ensure log integrity and identify log misbehavior. The lack of discussion may suggest that practitioners are either not yet engaging with these components or may not fully recognize their importance in correct use of the ecosystem. We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
5.1.4 Effects of Organization and Signing Tools on Identified Usability Concerns
        (a) Subjects from small organizations (6 subjects).         (b) Subjects from large organizations (7 subjects).
Figure 4: Tools vs. Evaluation Criteria. Reported usability issues across organizational contexts. Heatmaps show the number of unique participants raising each issue; darker cells indicate higher concentrations of reports. Abbreviations: DOC = Documentation (technical); TIME = Fulcio / Timestamping Workflow; INTEG = Integrations & Tooling; LOG = Log Monitoring Burden; PERF = Performance & Rate Limits; COST = Private Instance: Infrastructure & Cost; REG = Regulatory Suitability (Cross-factor); SUPP = Support & Maintenance; TLOG = Transparency Log Constraints; CMAT = Community Maturity/Support.
To provide additional context for the usability issues identified above, we highlight three observations from our analysis of organizational size and participant roles. See Figures˜4 and 9 for visualizations.
Large and small organizations share similar pain points
Across both large and small organizations, the densest cells occur for topics related to private/air-gapped deployments, transparency-log constraints, and documentation for customized/ private instances. In large organizations, technical leaders frequently surface additional concerns around support and maintenance expectations. In contrast, smaller organizations are more willing to combine Sigstore with other tooling to compensate for gaps. Notably, large organizations in our sample were more likely to retain internal/proprietary solutions or delay full Sigstore adoption. A usability lesson here is to prioritize private-deployment guidance (including monitoring playbooks), clarify support boundaries, and provide migration patterns and user stories for mixed stacks.
Scale effects: rate limits and latency concentrate in larger orgs
Performance and rate-limit issues are most prominent in medium/large organizations, consistent with higher CI/CD concurrency. This suggests the need for prescriptive CI recipes (e.g., batching, caching, retries/backoff, and concurrency guidance) alongside clearer Fulcio/OIDC and timestamping failure-mode documentation.
Medium organizations tend to “work around” with plain Sigstore
Participants from medium-sized organizations more often stick with plain Sigstore and work around gaps rather than self-host. This group benefits most from concise integration blueprints (GitHub/GitLab/Jenkins), ready-to-use attestation storage patterns, and targeted documentation updates, rather than heavy private-infrastructure guidance.
5.2 RQ2: Why Practitioners Change Tools
5.2.1 Drivers of Adoption Among Sigstore Users
@@
Most of our participants chose to switch to Sigstore due to poor experiences with previous tools and due to perceived advantages associated with Sigstore. We summarize these factors in Table˜11.
Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore. Some participants (e.g., P6) stated that their organizations have a policy of contributing to open-source communities they intend to adopt. Others noted that their organizations were relatively new compared to Sigstore—meaning that Sigstore had already existed before these organizations were established, and some of their members had contributed to Sigstore before joining their respective organizations. Additionally, some participants mentioned that key members of their organizations were directly involved in the initial development of the Sigstore project. P14 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community. There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.” While this pattern may partly reflect our interview sample, it highlights how prior involvement and community ties can strongly shape adoption decisions—a dynamic that could also apply more broadly in open-source ecosystems.
Available Sigstore Functionalities: Sigstore’s functionalities also influenced its adoption. While several participants were drawn to the transparency log (as highlighted in section˜5.1.1). {quoting} P5 (internal security tool, cloud security): “Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, [and] they also maintain a transparency log.”
@@
Regulations and Standards: The impact of regulations and standards on adopting Sigstore was significant among our participants. Some participants who previously used other tools (2) switched to Sigstore due to security frameworks and standards recommending it as a best practice. Additionally, the remaining participants (2) noted that regulations (e.g., E0-14028) played a role in their decision to adopt Sigstore.
@@
User Community & Trust Of Sigstore Creators: The large user community of Sigstore also motivated its adoption among participants. Conversely, participants P1 and P14 cited GPG’s smaller user base as a drawback. In addition, some participants expressed trust in industry consortia such as the CNCF (Cloud Native Computing Foundation), noting that Sigstore’s development and governance under CNCF gave it additional legitimacy and credibility compared to tools maintained by smaller, independent groups. {quoting} P3 (SAAS, SSC security tool): “We do defer often at the higher-level projects to different foundations. So we would be more likely to trust something from the CNCF…[than] independent developer projects.”
GPG & Notary Issues: Issues with previously existing tools played a major role in practitioners’ decisions to switch to Sigstore. Both GPG- and Notary-based signing implementations were reported as problematic. GPG has long been criticized for key management challenges and usability concerns, and our participants echoed these issues, citing additional factors such as low adoption rates, steep learning curves, and limited integration with modern technologies. Notary-related issues primarily concerned customer demands, compatibility with other technologies, and infrequent updates. Key and identity management challenges were common to both GPG and Notary. Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
5.2.2 Considerations from Non-Sigstore Users
Table 5: Non Sigstore Users: Reasons for Not Adopting Sigstore. This table summarizes the reasons non-users gave for not adopting Sigstore, grouped by their current signing tool and mapped to our usability framework theme.
@@ 5.2.2 Considerations from Non-Sigstore Users
In this section we highlight reasons mentioned by practitioners that have stopped them (or their organization) from switching to Next-gen tools like Sigstore. We also report issues experienced by these non-Sigstore users that have prompted a consideration (or in some cases partial adoption) of these next-gen tools.
@@
Barriers to Considering Sigstore: As highlighted in Table˜5, most reasons for non-adoption were organizational. {quoting} P8 (internal security tool & cloud APIs, internet services): “When it comes to open source projects and open source stuff, our organization is skeptical. Policy and everything drives teams away…There’s also the ongoing risk of having a product that’s built outside continually shipping code and being integrated into this product, where it deals with sensitive material. We do evaluate these things, but the scrutiny against code written outside of the company is extreme, and with good reason given the [supply chain] context.” Other notable issues arose from technical concerns, like privacy risks associated with Sigstore’s public instance and Macro environmental concern of the inherent risk of relying on a third-party tool maintained by open-source contributors.
@@
Factors Motivating Consideration for Sigstore: When practitioners were asked about issues with their current tools and potential fixes, most pointed to problems with existing tools as their main driver to alternatives, e.g., incorporating next-gen features into internal tools or partially adopting new solutions. Although some subjects reported drawing design inspiration from next-generation workflows, none specifically mentioned Sigstore. P11 (firmware & testing software, social technology) states, “I’m sure … [the internal signing tool] was influenced by [an open-source next-gen tool]…they adopted it and made the right tweaks internally.” Lack of transparency and integration of these tools to other platforms were also common. We summarize these factors in Table˜12.
5.2.3 Effects of Organizational Context
Figure 5: Push-pull-barrier heatmap.
Combining the results in section˜5.2.1 (Table˜11) and section˜5.2.2 (Tables˜12 and 5) with organizational size, former signing tools, and current signing tools reveals several patterns. We summarize these in the Push-pull-Barrier heatmap (Figure˜5)
Push, Pull, and Barrier factors influencing adoption
We grouped issues raised by participants into Push, Pull, and Barrier factors. We define these factors drawing on Bansal et al.’s Push-Pull-Mooring framework for consumer switching behavior [10]. Push factors are problems with prior tools that drove current users toward Sigstore. Pull factors are attractions of the new tool (and ecosystem). Barriers are reasons that stalled or prevented adoption.
While large organizations exhibit the greatest number of barriers, they report push and pull factors in roughly similar measure. By current-tool status, many push factors come from teams now using Sigstore alongside other mechanisms (e.g., attestations, internal tooling, PGP), which often reflects phased introduction of Sigstore into existing processes (and vice versa). A summary of the coded factors appears in Table˜13. A usability takeaway is that switching is primarily driven by human/social push (e.g., frustration with legacy workflows) and only secondarily by pull from Sigstore features and regulatory pressure (Sigstore being referenced by some standards). Thus, even as technical innovations “pull” adoption, community-building and other social supports also materially “push” change.
Non-adoption reasons are often organizational and non-technological
Non-adoption reasons are often organizational and non-technological. Although non-adopters are both pushed by shortcomings in internal tools and pulled by next-generation capabilities, adoption frequently stalls due to organizational constraints such as risk posture, support expectations, or regulatory requirements. This pattern is not unique to software signing—similar dynamics have been observed in other infrastructure domains (e.g., cloud technologies[6])—but it is particularly salient here because identity-based signing tools must be trusted across organizational boundaries. Our interpretation is therefore not that next-generation tools have failed, but rather that their technical attractiveness has created demand for more mature deployment models: private instances, enterprise-grade support, and stronger service guarantees.
Tool migration patterns
In our sample we observe recurring migration paths: PGP $\rightarrow$ Sigstore (sometimes retaining PGP during transition); Notary v1 $\rightarrow$ Sigstore or Notary v2; and internal/proprietary $\rightarrow$ Sigstore (or hybrid). Organizations mostly transition between tools slowly. Larger organizations are more likely to prefer internal/hybrid arrangements (reflecting policy, privacy, and support concerns), whereas small/medium organizations are more likely to adopt “pure” Sigstore sooner, using hybrids as transitional steps or to meet offline/private constraints.
6 Discussion
We highlight the main takeaways from our results, noting that this research is a case study of next-generation software signing tools (section˜2.1) using Sigstore. We discuss implications for this class of tools along three axes: a technological readiness assessment; their transformative potential; and the implications of the observed usability patterns
6.1 TRLs and Next-Generation Signing Tools
Because Sigstore, an exemplar of next-generation (identity-based) signing, is relatively recent, the maturity of its core capabilities is uneven. Understanding this maturity helps organizations assess implementation risk and prioritize investment [53]. Our usability study provides a preliminary signal. Drawing on participant reports and interpreting them through the Technology Readiness Level (TRL) lens [53], we note that TRL is a technology-maturation framework (not a usability scale). In practice, lower TRLs (concepts, prototypes) tend to surface mostly technical concerns, while higher TRLs (operational environments) bring non-technical constraints (organizational fit, governance, support/SLAs, regulation) to the foreground. Our results fit this progression: technological factors were most frequently discussed. Several capabilities, e.g., short-lived (ephemeral) keys, OIDC-based signer identity, and GitHub Actions workflows, received consistently positive mentions across different organizational contexts, suggesting higher maturity for these components (and analogous features in similar next-gen tools). In contrast, capabilities such as private/self-hosted deployments and operation in air-gapped or regulated environments surfaced moderate friction, indicating mid-level readiness and a need for stronger documentation, reference architectures, and support pathways. For organizations that have not yet adopted next-generation tools, these findings suggest comparatively lower risk for the first set of capabilities, while private/regulated deployments warrant careful evaluation and piloting. These assessments are preliminary and should be read in light of our case-study scope.
From Sigstore to the broader class: Sigstore demonstrates that identity-based signing can work at scale (across multiple organization sizes and use contexts in our sample), serving as a proof-of-concept for the class. Where we observed Sigstore-specific friction (e.g., private Fulcio/Rekor deployments, public-log privacy trade-offs), the underlying class implications are clear: enterprise deployment models, integrations, and operational guardrails are the next maturity hurdles for identity-based signing generally.
6.2 Transformative Potential
Sigstore, as the most widely deployed exemplar of next-generation (identity-based) signing, goes well beyond the minimal OIDC–key binding model offered by alternatives like OpenPubKey. In addition to identity-based signing and short-lived keys, Sigstore integrates a transparency log (Rekor) and broad ecosystem compatibility with attestations, container registries, and CI/CD workflows. Participants repeatedly emphasized these additional capabilities as unique differentiators, even while reporting usability frictions.
This dual reality—strong adoption coupled with persistent friction—underscores Sigstore’s role as a living testbed for both practice and research. Sigstore is already being leveraged in production ecosystems (e.g., Maven, PyPI, Kubernetes) and has catalyzed new research trajectories, such as privacy-preserving transparency [44] and diversified identity provisioning [57]. The broader implication is that Sigstore is not merely a tool to be adopted, but an infrastructure on which future signing solutions and empirical studies can (and are) being built. Widespread adoption makes its usability challenges consequential — and resolving them has cascading benefits across the supply chain. Its design decisions and shortcomings provide researchers with concrete leverage points for advancing the next generation of signing tools.
6.3 Lessons for Next-Gen Signing Usability
While our results reveal different adoption drivers between sigstore users and non users, there exists some shared usability concerns—between Sigstore users and non-users (section˜5.2). Both groups reported similar technological issues, particularly around integration with other systems and the transparency of signing infrastructures. This suggests cross-cutting concerns that affect not only Sigstore users but also non-users. A preliminary look at enhancement/issue trackers for other next-generation signing tools (e.g., OpenPubkey) shows similar themes [58]. Building on these observations, we distill actionable lessons for next-generation tools and outline opportunities for future work.